A simple pendulum is found to vibrate 50 times within 200s. When 1.5 meter of its length is reduced to a certain length it vibrates 50 times in 175s. Find the original length of the pendulum.

1
Expert's answer
2022-01-20T13:50:13-0500

Given:

"N_1=50,\quad t_1=200\:\rm s"

"N_2=50,\quad t_1=175\:\rm s"


The period of oscillation of a simple pendulum

"T_1=\frac{t_1}{N_1}=2\pi\sqrt{l/g}""T_2=\frac{t_2}{N_2}=2\pi\sqrt{(l-x)/g}"

"\left(\frac{t_1}{t_2}\right)^2=l/(l-x)""\left(\frac{200}{175}\right)^2=l/(l-1.5)""16\/9=l\/(l-1.5)"

"16l-24=9l"

"l=24/7=3.4\:\rm m"